OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3). y = 2x - 7 y = - one halfx - 2 y = 2x - 2 y = - one halfx - 7

OpenStudy (anonymous):

y = 2x - 7 so x= -7+2 which is -5 then you do -5 multiplied by 2 which is -10-7 which is -17. So y=-17

OpenStudy (anonymous):

so i say go with the first one:)

OpenStudy (anonymous):

thankkyaa(:

OpenStudy (anonymous):

no problem:)